OpenStudy (anonymous):

What is the length of the focal width? x=(1/4)y^2

OpenStudy (dumbcow):

=1/4a where a = 1/4

OpenStudy (anonymous):

so the answer is 1/4?

OpenStudy (dumbcow):

substitute 1/4 for a \[\frac{1}{4*(1/4)} = 1\] focus at (1,0)

OpenStudy (anonymous):

thanks!